Problem Page

Gaseous methane CH4 will react with gaseous oxygen O2 to produce gaseous carbon dioxide CO2 and gaseous water H2O. Suppose 10. g of methane is mixed with 22.6 g of oxygen. Calculate the maximum mass of water that could be produced by the chemical reaction. Round your answer to 3 significant digits.
Answer #1

Molar mass of CH4,

MM = 1*MM(C) + 4*MM(H)

= 1*12.01 + 4*1.008

= 16.042 g/mol

mass(CH4)= 10.0 g

use:

number of mol of CH4,

n = mass of CH4/molar mass of CH4

=(10 g)/(16.04 g/mol)

= 0.6234 mol

Molar mass of O2 = 32 g/mol

mass(O2)= 22.6 g

use:

number of mol of O2,

n = mass of O2/molar mass of O2

=(22.6 g)/(32 g/mol)

= 0.7063 mol

Balanced chemical equation is:

CH4 + 2 O2 ---> 2 H2O + CO2

1 mol of CH4 reacts with 2 mol of O2

for 0.6234 mol of CH4, 1.247 mol of O2 is required

But we have 0.7063 mol of O2

so, O2 is limiting reagent

we will use O2 in further calculation

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

According to balanced equation

mol of H2O formed = (2/2)* moles of O2

= (2/2)*0.7063

= 0.7063 mol

use:

mass of H2O = number of mol * molar mass

= 0.7063*18.02

= 12.72 g

Answer: 12.7 g